) Determine whether each of these functions is a bijection from R to R.
(i) f(x) = -3x + 4
(ii) f(x) = 2x + 1
(iii) f(x) = x
2 + 1
(iv) f(x) = -3x
2 + 7
(v) f(x) = (x+1)
(x+2)
1
Expert's answer
2020-09-03T16:23:22-0400

(i) "\forall x \in R\; \; \exist !y = -3x+4" and "\forall y \in R \; \exist!x= \frac{4-y}{3} \Rightarrow" bijection

(ii) "\forall x \in R\; \; \exist !y = 2x+1" and "\forall y \in R \; \exist!x= \frac{y-1}{2} \Rightarrow" bijection

(iii) "y = x^2+1", "y(-1) = y(1) = 2 \Rightarrow" is not injective, thus not bijection

(iv) "y=-3x^2+7" , "y(-1) = y(1) = 4 \Rightarrow" is not injective, thus not bijection

(v) "y = (x+1)(x+2)= x^2+2x+2 = (x+1)^2+1", "y(-3) = y(1) = 5 \Rightarrow" is not injective, thus not bijection